OpenStudy (anonymous):

A store displays six computers on a shelf side–by–side. If the first computer is seven inches wide and each successive computer is two inches wider than the previous one, find the total width of the computers on the shelf. Part 1: Describe the sigma notation used in answering the question above. (2 points) Part 2: Show all your work and answer the question. (3 points)

OpenStudy (anonymous):

so how do i solve that?

OpenStudy (ghazi):

just put n=1 for 2 computers or as your question says you have 6 computers put n=5 you'll have width = 7+2*5=17

OpenStudy (ghazi):

no wait

OpenStudy (anonymous):

dont i have to use that symbol at some point..?

OpenStudy (ghazi):

it will have series of 7*n+2*x where n=number of computers and x= n-1

OpenStudy (anonymous):

i am sorry i dont know how to slove that can you solve it for me?

OpenStudy (ghazi):

put n=6 and x=5 your total width will be = 7*6+2*5= 42+10=52 inches

OpenStudy (anonymous):

thank you!!!!

OpenStudy (ghazi):

:)